OpenStudy (maheshmeghwal9):

Aftab tells his daughter,"Seven years ago,I was seven times old as you were then.Also,three years from now,I shall be three times as old as you will be."Represent this situation algebraically & graphically.

OpenStudy (maheshmeghwal9):

@eSpeX

OpenStudy (maheshmeghwal9):

Plz help!

OpenStudy (anonymous):

Let Aftab's present age be x and his daughter's present age be y. Seven years ago, Aftab's age = x-7 Daughter's age = y-7 By the given condition, x-7 = 7 *(y-7) Three years hence, Aftab's age = x+3 Daughter's age = y+3 By the given condition, x+3 = 3*(y+3) x - 7 = 7y - 49 x + 3 = 3y + 9 On solving both the equations, you will arrive at the required answer.

OpenStudy (maheshmeghwal9):

K!but how to represent this situation graphically?@Spacemanifestation

OpenStudy (espex):

To represent the problem graphically, you could produce a bar graph, line graph, or even a scatter chart. You could set Aftab's age as the vertical axis and his daughters age as the horizontal axis and plot the relationship between the two ages.